• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ibrahim

    Log entry time 08:04:41 on February18,2009

    Entry number 261797

    keyword=Owl Shift Summary (Feb. 18, 2009)


    Owl Shift Summary (Feb. 18, 2009)

    00:00 Shift Started. Continue runs 1531/20239 (production, half wave plate is OUT, 15uA).

    00:18 Run 1532 (5M)

    00:53 Run 1533

    01:53 Run 1534

    02:15 Change the half wave plate from OUT to IN.

    02:25 Run 1535 (half wave plate is IN).

    03:09 NMR frequency sweep measurement (59.58%).

    03:18 Runs 1537/20240 (run 1536 is junk).

    03:58 Run 1538

    04:31 Run 1539

    05:00 The Compton vncviewer crashed in the back room (Rect too big!). Had to restart to get access to the Compton DAQ controls again. The new runs are:

    Compton 15822
    FADC 60563
    HAPPEX 30817

    05:07 Run 1540

    05:43 Run 1541 (2M only)

    06:00 Called MCC to change the half wave plate from IN to OUT). Restarted the Compton and HAPPEX DAQs (back room) for the half wave plate change. The new runs are:

    Compton 15823
    FADC 60564
    HAPPEX 30818

    06:05 Run 1542 (half wave plate is OUT).

    06:48 Run 1543

    06:54 LHRS Run 20242 (run 20241 is Junk).

    07:00 The Compton vncviewer crashed again (same error: Rect too big!). Restarted again.

    07:24 Run 1544

    08:00 Shift Ends